# Generalised Flow Maps for Few-Step Generative Modelling on Riemannian Manifolds

### v1

@@ 1 Introduction
Dynamical measure transport offers a unifying and prescriptive framework for constructing neural network-based generative models that learn to sample a desired target distribution by pushing forward a tractable prior.Numerical solution of the resulting dynamical systems has led to popular method families—including diffusion models (Song & Ermon, 2020), flow matching (Lipman et al., 2022; Peluchetti, 2023; Liu et al., 2022; Albergo & Vanden-Eijnden, 2022), and stochastic interpolants (Albergo et al., 2023)—which together have revolutionised the field and led to state-of-the-art results over continuous modalities (Karras et al., 2024; Polyak et al., 2024).While often applied to Euclidean data such as images, this powerful paradigm naturally extends to data types that are inherently geometric and lie on a known Riemannian manifold, in which case the associated flows and diffusions are defined directly on the manifold (Huang et al., 2022b; De Bortoli et al., 2022; Chen & Lipman, 2024).Such geometric generative models have found widespread application in high-impact scientific settings such as rational drug design on the $\mathrm{SE(3)}^{\scriptscriptstyle N}$ manifold of protein backbones (Bose et al., 2024; Huguet et al., 2024; Watson et al., 2023), generative material design in computational chemistry (Miller et al., 2024), and even discrete data using the Fisher-Rao geometry on the probability simplex (Davis et al., 2024; Cheng et al., 2025a).The scalability of dynamical transport-based generative models arises from their use of simple regression-based objectives that lead to rapid simulation-free training.However, unlike training, obtaining high-quality generated samples at inference time requires numerical solution of the parametrised dynamical system, which necessitates evaluating the large learnt model numerous times.The computational complexity of inference is further burdened in geometric settings, where each step of simulation requires computing potentially numerically unstable operators that must guarantee manifold constraints for faithful inference.For instance, for matrix Lie groups like $\mathrm{SO(3)}$, computing the standard exponential and logarithmic maps requires truncating an infinite matrix power series (Al-Mohy & Higham, 2012), which may incur additional sources of error beyond just step-size discretisation error compared to inference on Euclidean spaces. Therefore, reducing the number of Riemannian operations, while maintaining this powerful inductive bias, has the potential to lead to higher fidelity generated samples.
@@
The search for accelerated inference with dynamical generative models defined on Euclidean spaces has spurred the development of fast inference techniques that fall broadly into two methodological families.The first family involves the intricate design of higher-order samplers (Dockhorn et al., 2022a; Zhang & Chen, 2022; Sabour et al., 2024; Karras et al., 2022), which can produce high-quality samples with less than ten steps.The second pursues direct learning of the flow map—the global solution of the governing differential equation—rather than the instantaneous vector field used in standard simulation (Boffi et al., 2024).This flow map perspective underlies distillation-based approaches, the most popular of which include consistency models (Song et al., 2023b; Song & Dhariwal, 2023; Lu & Song, 2024), which attempt to learn the one-step flow map of a pre-trained teacher.Distillation-based methods have also been extended to the multi-step flow map setting we consider here at scale (Sabour et al., 2025).Recent advances have further demonstrated that state-of-the-art one-step generative models can be trained from scratch without a pre-trained teacher through self-consistent self-distillation (Boffi et al., 2025; Frans et al., 2025) or Monte Carlo direct training (Geng et al., 2025) objectives.These exciting developments naturally raise the question:
@@ 2.2 Stochastic Interpolants and Flow Maps
Setting $\alpha_{t}\equiv 1-t$ and $\beta_{t}\equiv t$ recovers well known methods such as flow-matching (Albergo & Vanden-Eijnden, 2022; Lipman et al., 2022) and rectified flow (Liu et al., 2022), while imposing optimal transport costs on the coupling results in OT-CFM (Tong et al., 2023).Flow maps in $\mathbb{R}^{d}$. An alternative path to generating samples is to instead learn the integrator of the probability flow ODE directly, in order to avoid, at inference time, the costly numerical integration. Introduced in Boffi et al. (2024), flow maps are functions $X:[0,1]^{2}\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ that satisfy the jump condition: $X_{s,t}(x_{s})=x_{t}$, where $(x_{t})_{t\in[0,1]}$ is any solution of the probability flow. One can thus sample from $\rho_{1}$ by first sampling $x_{0}\sim\rho_{0}$ and then applying $X_{0,1}(x_{0})\sim\rho_{1}$. These are parametrised as $X^{\theta}_{s,t}(x_{s})=x+(t-s)v^{\theta}_{s,t}(x_{s})$, for any $x$, $s$ and $t$, so that the boundary condition, at $t=s$, $X^{\theta}_{t,t}=\mathrm{Id}$ is automatically satisfied. There are various self-distillation (i.e., from scratch) objectives for learning a flow-map, which are used in conjunction with typical flow-matching objectives (Albergo et al., 2023; Tong et al., 2023; Lipman et al., 2023),For brevity, the three (Euclidean) flow map self-distillation loss variations—Eulerian, Lagrangian, and progressive—are provided in §A.
@@ 3.1 Generalised Flow Maps
The proof for Lemma 1 is provided in §B.2 and also illustrated in Figure 1. The lemma underscores the key idea that, for $s=t$ (on the diagonal of the times space, $0,1]^{2}$), the derivative of the GFM is the instantaneous vector field, $v_{t}$ of equation 4.Moreover, from the parametrisation in Proposition 1, and from the fact that $\lim_{s\to t}\partial_{t}X_{s,t}(x)=v_{t,t}(x)$ for any $x$, it follows that $v_{t,t}=v_{t}$ for any $t\in[0,1]$. (We thoroughly prove the limit in Lemma 4.)Therefore, we hereinafter use $v_{t,t}$ and $v_{t}$ interchangeably, and we may train the GFM on the diagonal $s=t$ using Riemannian Flow Matching (RFM) (Chen & Lipman, 2024), the loss of which is given by:
@@ 4 Experiments
We test the empirical caliber of GFM on a suite of standard geometric generative modelling benchmarks.  [^1] We make our code publicly available at the following URL: https://github.com/olsdavis/gfm. Specifically, we instantiate GFM on torsion angles ($\mathbb{T}^{2}\cong\mathbb{S}^{1}\times\mathbb{S}^{1}$) found in protein side chains and RNA backbones $\mathbb{T}^{7}\cong(\mathbb{S}^{1})^{7}$ (Lovell et al., 2003; Murray et al., 2003), catastrophic geospatial events on Earth ($\mathbb{S}^{2}$) as introduced in Mathieu & Nickel (2020a), a synthetic dataset of the manifold of $3\text{D}$ rotations ($\mathrm{SO(3)}$), and on the Poincaré disk (hyperbolic geometry).Metrics.To evaluate the few-step sample quality of our models, we use the empirical MMD (Maximum Mean Discrepancy) between the test-set and the samples, with an RBF kernel using the manifold’s distance, $d_{g}$, and a bandwidth of $\kappa=1$. See §C for the calculation details. To assess the learnt vector field, we also compute the negative log-likelihood (NLL) on the test set when available. Finally, we provide qualitative samples to further assert the method’s soundness.Baselines. We use RFM (Chen & Lipman, 2024), which is the state-of-the-art method, as the main baseline. Additionally, we include results for a Riemannian diffusion model (RDM)  (Huang et al., 2022b) and Riemannian score-based generative models (RSGM) (De Bortoli et al., 2022) with test NLL results taken directly from the respective papers due to a lack of open source code. We also report a mixture of power spherical distributions (MoPS) (De Cao & Aziz, 2020) for test NLL.
@@ 4.1 Proteins Torsion Angles and RNA backbones on flat tori
RFM (Chen & Lipman, 2024) ${1.01}\pm 0.025$ $\mathbf{1.90\pm 0.055}$ $0.15\pm 0.027$ $1.18\pm 0.055$ $\mathbf{-5.20\pm 0.067}$
@@ 4.2 Earth catastrophes on the 2-sphere
We evaluate GFM on a collection of geospatial data, represented on the 2-sphere, $\mathbb{S}^{2}$. The dataset was first introduced in Mathieu & Nickel (2020b) and is curated from various sources (NOAA, 2020a; b; Brakenridge, 2017; EOSDIS, 2020). We report our results in Figures 4 and 3. Analogous to tori, we observe that the implicit flow $v^{\theta}_{t,t}$ within GFM offers log-likelihoods that outperform those of RFM and all other methods on three out of four datasets, with the exception of “Volcanoes”. To assess sample quality, we again plot the MMD as a function of the number of integration steps, and observe once more great improvement on MMD for all our methods at 1 or 2 NFEs, while preserving high quality for higher NFEs. Finally, we provide the densities learnt by RFM and G-LSD on 2D Earth plots in Figure 5 (with the remaining plots in Figure 7), and remark similar likelihood landscapes between RFM and G-LSD.
@@
Riemannian CNF  (Mathieu & Nickel, 2020a) $0.97{\pm 0.15}$ $0.19{\pm 0.04}$ $0.90{\pm 0.03}$ $0.66{\pm 0.05}$
@@
RFM (Chen & Lipman, 2024) $\mathbf{-7.93\pm 1.67}$ $0.28\pm 0.08$ $0.42\pm 0.05$ $1.86\pm 0.11$
@@ 5 Related work
Riemannian generative models. The most related early efforts to build manifold structure into generative models come from conventional normalising flows built out of iterative coupling layers (Tabak & Vanden-Eijnden, 2010; Rezende & Mohamed, 2016; Dinh et al., 2017), where each coupling layer was designed to preserve the manifold structure (Bose et al., 2020; Rezende et al., 2020; Kanwar et al., 2020; Boyda et al., 2021; Bose et al., 2021). These ideas have also been extended to continuous-time flows and diffusions on general Riemannian structures (Lou et al., 2020; Mathieu & Nickel, 2020a; Falorsi & Forré, 2020; Bortoli et al., 2022), and their optimisation has also been made simulation-free (Rozen et al., 2021; Chen & Lipman, 2024; Bose et al., 2020; Davis et al., 2024). Here, our focus is on extending these approaches so that the more efficient any-step flow map is well-posed and learnable on general geometries.Accelerated inference in generative models. Early work on accelerated inference focused on a teacher-student procedure (Song et al., 2020; Luhman & Luhman, 2021; Salimans & Ho, 2022; Meng et al., 2023) where an expensive inference model is distilled to produce the same output in fewer steps. Orthogonal to this have been efforts to parallel diffusion inference with adaptive (Chen et al., 2024; Dockhorn et al., 2022b; Shih et al., 2023; Tang et al., 2024) or speculative (Bortoli et al., 2025) schemes. Consistency models aim to directly learn the one-step map to the data distribution from any point along the trajectory (Song et al., 2023b; Kim et al., 2024; Song & Dhariwal, 2023). The flow map (Boffi et al., 2024) has emerged as a unifying picture, and recent efforts have shown how to distill it (Sabour et al., 2024) or directly learn it (Boffi et al., 2025; Geng et al., 2025) using the equations that characterise it. We take the leap here to show how to generalise this complete class of models to the Riemannian setting for performant ends.Concurrent work. Concurrent and most related to our work is that of Cheng et al. (2025b). The authors propose a new method, Riemannian Consistency Models (RCM), to train few-step generative models on Riemannian manifolds from scratch. Their approach directly ports the original Consistency Models (Song et al., 2023b) to Riemannian geometries, relying on more sophisticated geometric constructions. In contrast, GFM enjoys simpler practical instantiations as it relies on the self-distillation of the Flow Maps framework (Boffi et al., 2025) and recovers shortcut models as a special case of the PSD objective executed on Riemannian manifolds.
@@ Appendix A Self-Distillation on Euclidean spaces
    $\partial_{s}X_{s,t}(x)+\nabla X_{s,t}(x)b_{s}(x)=0\qquad\forall\>(x,s,t)\in\mathbb{R}^{d}\times[0,1]^{2}.$ (17)
@@
The recommendations in equation 31 and equation 33 are precisely what would arise given a pre-trained teacher.The recommendation in equation 32 is similar, but also places a $\mathrm{stopgrad}(\cdot)$ on the spatial Jacobian of the model, which often improves optimization stability significantly.Mean Flows and consistency training.Given the choice of $\mathrm{stopgrad}(\cdot)$ in equation 32, the resulting parameter gradient will be linear in $v^{\theta}_{s,s}$.We may then replace $v^{\theta}_{s,s}(I_{s})$ by the Monte Carlo estimate $\partial_{t}{I}_{s}$ of the ideal flow $b_{s}(x)=\mathbb{E}[\partial_{t}{I}_{s}|I_{s}=x]$, using the tower property of the conditional expectation $\mathbb{E}\left[f(I_{s})\partial_{t}{I}_{s}\right]=\mathbb{E}\left[f(I_{s})\mathbb{E}\left[\partial_{t}{I}_{s}|I_{s}\right]\right]=\mathbb{E}\left[f(I_{s})v_{s}(I_{s})\right]$ for any function $f:\mathbb{R}^{d}\to\mathbb{R}^{d}$.Without this choice of $\mathrm{stopgrad}$, this replacement is not possible due to the quadratic term $\mathbb{E}[|\nabla X^{\theta}_{s,t}(I_{s})\partial_{t}{I}_{s}|^{2}]\neq\mathbb{E}[|\nabla X^{\theta}_{s,t}(I_{s})v_{s}(I_{s})|^{2}]$ because of the nonlinearity in $|\cdot|^{2}$.With this $\mathrm{stopgrad}$, the quadratic term is a constant from the perspective of the gradient and hence this discrepancy can be ignored.This trick is used by both consistency training (Song et al., 2023a; Lu & Song, 2025) and Mean Flows (Geng et al., 2025); in fact, these algorithms can be recovered by expanding

### v2

@@ 1 Introduction
Dynamical measure transport offers a unifying and prescriptive framework for constructing neural network-based generative models that learn to sample a desired target distribution by pushing forward a tractable prior.Numerical solution of the resulting dynamical systems has led to popular method families—including diffusion models (Song and Ermon, 2020), flow matching (Lipman et al., 2022; Peluchetti, 2023; Liu et al., 2022; Albergo and Vanden-Eijnden, 2022), and stochastic interpolants (Albergo et al., 2023)—which together have revolutionised the field and led to state-of-the-art results over continuous modalities (Karras et al., 2024; Polyak et al., 2024).While often applied to Euclidean data such as images, this powerful paradigm naturally extends to data types that are inherently geometric and lie on a known Riemannian manifold, in which case the associated flows and diffusions are defined directly on the manifold (Huang et al., 2022b; De Bortoli et al., 2022; Chen and Lipman, 2024).Such geometric generative models have found widespread application in high-impact scientific settings such as rational drug design on the $\mathrm{SE(3)}^{\scriptscriptstyle N}$ manifold of protein backbones (Bose et al., 2024; Huguet et al., 2024; Watson et al., 2023), generative material design in computational chemistry (Miller et al., 2024), and even discrete data using the Fisher-Rao geometry on the probability simplex (Davis et al., 2024; Cheng et al., 2025a).The scalability of dynamical transport-based generative models arises from their use of simple regression-based objectives that lead to rapid simulation-free training.However, unlike training, obtaining high-quality generated samples at inference time requires numerical solution of the parametrised dynamical system, which necessitates evaluating the large learnt model numerous times.The computational complexity of inference is further burdened in geometric settings, where each step of simulation requires computing potentially numerically unstable operators that must guarantee manifold constraints for faithful inference.For instance, for matrix Lie groups like $\mathrm{SO(3)}$, computing the standard exponential and logarithmic maps requires truncating an infinite matrix power series (Al-Mohy and Higham, 2012), which may incur additional sources of error beyond just step-size discretisation error compared to inference on Euclidean spaces. Therefore, reducing the number of Riemannian operations, while maintaining this powerful inductive bias, has the potential to lead to higher fidelity generated samples.
@@
The search for accelerated inference with dynamical generative models defined on Euclidean spaces has spurred the development of fast inference techniques that fall broadly into two methodological families.The first family involves the intricate design of higher-order samplers (Dockhorn et al., 2022a; Zhang and Chen, 2022; Sabour et al., 2024; Karras et al., 2022), which can produce high-quality samples with less than ten steps.The second pursues direct learning of the flow map—the global solution of the governing differential equation—rather than the instantaneous vector field used in standard simulation (Boffi et al., 2024).This flow map perspective underlies distillation-based approaches, the most popular of which include consistency models (Song et al., 2023b; Song and Dhariwal, 2023; Lu and Song, 2024), which attempt to learn the one-step flow map of a pre-trained teacher.Distillation-based methods have also been extended to the multi-step flow map setting we consider here at scale (Sabour et al., 2025).Recent advances have further demonstrated that state-of-the-art one-step generative models can be trained from scratch without a pre-trained teacher through self-consistent self-distillation (Boffi et al., 2025; Frans et al., 2025) or Monte Carlo direct training (Geng et al., 2025) objectives.These exciting developments naturally raise the question:
@@ 2.2 Stochastic Interpolants and Flow Maps
Setting $\alpha_{t}\equiv 1-t$ and $\beta_{t}\equiv t$ recovers well known methods such as flow-matching (Albergo and Vanden-Eijnden, 2022; Lipman et al., 2022) and rectified flow (Liu et al., 2022), while imposing optimal transport costs on the coupling results in OT-CFM (Tong et al., 2023).Flow maps in $\mathbb{R}^{d}$. An alternative path to generating samples is to instead learn the integrator of the probability flow ODE directly, in order to avoid, at inference time, the costly numerical integration. Introduced in Boffi et al. (2024), flow maps are functions $X:[0,1]^{2}\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ that satisfy the jump condition: $X_{s,t}(x_{s})=x_{t}$, where $(x_{t})_{t\in[0,1]}$ is any solution of the probability flow. One can thus sample from $\rho_{1}$ by first sampling $x_{0}\sim\rho_{0}$ and then applying $X_{0,1}(x_{0})\sim\rho_{1}$. These are parametrised as $X^{\theta}_{s,t}(x_{s})=x+(t-s)v^{\theta}_{s,t}(x_{s})$, for any $x$, $s$ and $t$, so that the boundary condition, at $t=s$, $X^{\theta}_{t,t}=\mathrm{Id}$ is automatically satisfied. There are various self-distillation (i.e., from scratch) objectives for learning a flow-map, which are used in conjunction with typical flow-matching objectives (Albergo et al., 2023; Tong et al., 2023; Lipman et al., 2023),For brevity, the three (Euclidean) flow map self-distillation loss variations—Eulerian, Lagrangian, and progressive—are provided in §A.
@@ 3.1 Generalised Flow Maps
The proof for Lemma 1 is provided in §B.2 and also illustrated in Figure 1. The lemma underscores the key idea that, for $s=t$ (on the diagonal of the times space, $0,1]^{2}$), the derivative of the GFM is the instantaneous vector field, $v_{t}$ of equation 4.Moreover, from the parametrisation in Proposition 1, and from the fact that $\lim_{s\to t}\partial_{t}X_{s,t}(x)=v_{t,t}(x)$ for any $x$, it follows that $v_{t,t}=v_{t}$ for any $t\in[0,1]$. (We thoroughly prove the limit in Lemma 4.)Therefore, we hereinafter use $v_{t,t}$ and $v_{t}$ interchangeably, and we may train the GFM on the diagonal $s=t$ using Riemannian Flow Matching (RFM) (Chen and Lipman, 2024), the loss of which is given by:
@@ 4 Experiments
We test the empirical calibre of GFM on a suite of standard geometric generative modelling benchmarks. Specifically, we instantiate GFM on torsion angles ($\mathbb{T}^{2}\cong\mathbb{S}^{1}\times\mathbb{S}^{1}$) found in protein side chains and RNA backbones $\mathbb{T}^{7}\cong(\mathbb{S}^{1})^{7}$ (Lovell et al., 2003; Murray et al., 2003), catastrophic geospatial events on Earth ($\mathbb{S}^{2}$) as introduced in Mathieu and Nickel (2020a), a synthetic dataset of the manifold of $3\text{D}$ rotations ($\mathrm{SO(3)}$), and on the Poincaré disk (hyperbolic geometry).Metrics.To evaluate the few-step sample quality of our models, we use the empirical MMD (Maximum Mean Discrepancy) between the test-set and the samples, with an RBF kernel using the manifold’s distance, $d_{g}$, and a bandwidth of $\kappa=1$. See §C for the calculation details. To assess the learnt vector field, we also compute the negative log-likelihood (NLL) on the test set when available. We have included in Section C.1 a discussion explaining why NLL is relevant in this context. Finally, we provide qualitative samples to further assert the method’s soundness.Baselines. We use RFM (Chen and Lipman, 2024), which is the state-of-the-art method, as the main baseline. Additionally, we include results for a Riemannian diffusion model (RDM)  (Huang et al., 2022b) and Riemannian score-based generative models (RSGM) (De Bortoli et al., 2022) with test NLL results taken directly from the respective papers due to a lack of open source code. We also report a mixture of power spherical distributions (MoPS) (De Cao and Aziz, 2020) for test NLL.Finally, we also detailed the computational cost of having using our proposed methods against RFM in Section C.4.
@@ 4.1 Proteins Torsion Angles and RNA backbones on flat tori
RFM (Chen and Lipman, 2024) ${1.01}\pm 0.025$ $\mathbf{1.90\pm 0.055}$ $0.15\pm 0.027$ $1.18\pm 0.055$ $\mathbf{-5.20\pm 0.067}$
@@ 4.2 Earth catastrophes on the 2-sphere
We evaluate GFM on a collection of geospatial data, represented on the 2-sphere, $\mathbb{S}^{2}$. The dataset was first introduced in Mathieu and Nickel (2020b) and is curated from various sources (NOAA, 2020a; b; Brakenridge, 2017; EOSDIS, 2020). We report our results in Figures 4 and 3. Analogous to tori, we observe that the implicit flow $v^{\theta}_{t,t}$ within GFM offers log-likelihoods that outperform those of RFM and all other methods on three out of four datasets, with the exception of “Volcanoes”. To assess sample quality, we again plot the MMD as a function of the number of integration steps, and observe once more great improvement on MMD for all our methods at 1 or 2 NFEs, while preserving high quality for higher NFEs. Finally, we provide the densities learnt by RFM and G-LSD on 2D Earth plots in Figure 5 (with the remaining plots in Figure 7), and remark similar likelihood landscapes between RFM and G-LSD.
@@
Riemannian CNF  (Mathieu and Nickel, 2020a) $0.97{\pm 0.15}$ $0.19{\pm 0.04}$ $0.90{\pm 0.03}$ $0.66{\pm 0.05}$
@@
RFM (Chen and Lipman, 2024) $\mathbf{-7.93\pm 1.67}$ $0.28\pm 0.08$ $0.42\pm 0.05$ $1.86\pm 0.11$
@@ 5 Related work
Riemannian generative models. The most related early efforts to build manifold structure into generative models come from conventional normalising flows built out of iterative coupling layers (Tabak and Vanden-Eijnden, 2010; Rezende and Mohamed, 2016; Dinh et al., 2017), where each coupling layer was designed to preserve the manifold structure (Bose et al., 2020; Rezende et al., 2020; Kanwar et al., 2020; Boyda et al., 2021; Bose et al., 2021). These ideas have also been extended to continuous-time flows and diffusions on general Riemannian structures (Lou et al., 2020; Mathieu and Nickel, 2020a; Falorsi and Forré, 2020; Bortoli et al., 2022), and their optimisation has also been made simulation-free (Rozen et al., 2021; Chen and Lipman, 2024; Bose et al., 2020; Davis et al., 2024). Here, our focus is on extending these approaches so that the more efficient any-step flow map is well-posed and learnable on general geometries.Accelerated inference in generative models. Early work on accelerated inference focused on a teacher-student procedure (Song et al., 2020a; Luhman and Luhman, 2021; Salimans and Ho, 2022; Meng et al., 2023) where an expensive inference model is distilled to produce the same output in fewer steps. Orthogonal to this have been efforts to parallel diffusion inference with adaptive (Chen et al., 2024; Dockhorn et al., 2022b; Shih et al., 2023; Tang et al., 2024) or speculative (Bortoli et al., 2025) schemes. Consistency models aim to directly learn the one-step map to the data distribution from any point along the trajectory (Song et al., 2023b; Kim et al., 2024; Song and Dhariwal, 2023). The flow map (Boffi et al., 2024) has emerged as a unifying picture, and recent efforts have shown how to distill it (Sabour et al., 2024) or directly learn it (Boffi et al., 2025; Geng et al., 2025) using the equations that characterise it. We take the leap here to show how to generalise this complete class of models to the Riemannian setting for performant ends.Concurrent work. Concurrent and most related to our work is that of Cheng et al. (2025b). The authors propose a new method, Riemannian Consistency Models (RCM), to train few-step generative models on Riemannian manifolds from scratch. Their approach directly ports the original Consistency Models (Song et al., 2023b) to Riemannian geometries, relying on more sophisticated geometric constructions. In contrast, GFM enjoys simpler practical instantiations as it relies on the self-distillation of the Flow Maps framework (Boffi et al., 2025) and recovers shortcut models as a special case of the PSD objective executed on Riemannian manifolds.
@@ Acknowledgements
Ethics Statement
We hereby acknowledge and declare to have abided by the ICLR Code of Ethics. Our current work does not utilise any sensitive data, nor does it directly enable nefarious usage, although other derivative works could use ours in unethical contexts. The authors declare no competing financial interests.
Reproducibility Statement
All of our results are reproducible and have been run with the same 5 random seeds that were set in advance, and that were never changed throughout our evaluation, to provide us our standard deviations. The datasets we use are entirely public and are freely available to all. Our code will be made public upon acceptance alongside detailed instructions for running it locally, and the exact configuration files that were used.
@@ Appendix A Self-Distillation on Euclidean spaces
    $\partial_{s}X_{s,t}(x)+\nabla X_{s,t}(x)v_{s}(x)=0\qquad\forall\>(x,s,t)\in\mathbb{R}^{d}\times[0,1]^{2}.$ (17)
@@
The recommendations in equation 31 and equation 33 are precisely what would arise given a pre-trained teacher.The recommendation in equation 32 is similar, but also places a $\mathrm{stopgrad}(\cdot)$ on the spatial Jacobian of the model, which often improves optimization stability significantly.Mean Flows and consistency training.Given the choice of $\mathrm{stopgrad}(\cdot)$ in equation 32, the resulting parameter gradient will be linear in $v^{\theta}_{s,s}$.We may then replace $v^{\theta}_{s,s}(I_{s})$ by the Monte Carlo estimate $\partial_{t}{I}_{s}$ of the ideal flow $b_{s}(x)=\mathbb{E}[\partial_{t}{I}_{s}|I_{s}=x]$, using the tower property of the conditional expectation $\mathbb{E}\left[f(I_{s})\partial_{t}{I}_{s}\right]=\mathbb{E}\left[f(I_{s})\mathbb{E}\left[\partial_{t}{I}_{s}|I_{s}\right]\right]=\mathbb{E}\left[f(I_{s})v_{s}(I_{s})\right]$ for any function $f:\mathbb{R}^{d}\to\mathbb{R}^{d}$.Without this choice of $\mathrm{stopgrad}$, this replacement is not possible due to the quadratic term $\mathbb{E}[|\nabla X^{\theta}_{s,t}(I_{s})\partial_{t}{I}_{s}|^{2}]\neq\mathbb{E}[|\nabla X^{\theta}_{s,t}(I_{s})v_{s}(I_{s})|^{2}]$ because of the nonlinearity in $|\cdot|^{2}$.With this $\mathrm{stopgrad}$, the quadratic term is a constant from the perspective of the gradient and hence this discrepancy can be ignored.This trick is used by both consistency training (Song et al., 2023a; Lu and Song, 2025) and Mean Flows (Geng et al., 2025); in fact, these algorithms can be recovered by expanding